You are here: CLASSE Wiki>ACC/Bunch Web>BSM23ELog>BSM23ELog_2007 (18 Jun 2007, EugeneTanke)Edit Attach
Tags

BSM23E Vertical Beamsize Monitor Electronic Logbook 2007

03 Oct 2022 - 18:11

January 2007

January 2, 2007 (C.Strohman, T.Wilksen, E. Tanke): Upgrade of Xlinx and Coldfire software

  • The new version of the Xilinx software now running on this unit is 5.5
  • The new version of the Coldfire software now running on this unit is 0.96B

January 4, 2007 (R.Holtzapple, E. Tanke): CESR-c beamsize measurements

  • Present CESR-C conditions are that we have current in the following 23 bunches:
    • T1B4
    • T2B3,T2B4,T2B5
    • T3B3,T3B4,T3B5
    • T4B3,T4B4,T4B5
    • T5B3,T5B4,T5B5
    • T6B3,T6B4,T6B5
    • T7B3,T7B4,T7B5
    • T8B3,T8B4,T8B5
    • T9B5
  • A typical run is currently ~ 30 minutes. 10k turn (raw) measurements using T1B3 as pedestal bunch:
    • Beginning run "A" (at 13:28): BSM23W datafile 936, BSM23E datafile 935
    • End run "A" (at 14:00): BSM23W datafile 937
    • Different end of run (at 15:14:): BSM23W datafile 939, BSM23E datafile 938
    • Files are currently on the VAX : [cesr.palmer.ts101.bsm]

January 5, 2007 (E. Tanke): ADC buffer read-out problems

  • Upon testing the ADC buffer read-out via the ethernet, it was found that same memory locations held data different from those when read out via the Xbus
  • Upon consecutive read-outs data via the ethernet, data was found to be different for certain memory locations
  • Also when making consecutive read-outs via the Xbus, data was found to be different for certain memory locations
  • In order to further investigate this problem, this DSP unit will be exchanged for the on in the DAQlab coming Tuesday
  • Seen the above, it is most likely that yesterdays' measurements are to some extent corrupted

January 9, 2007 (E. Tanke): Swapped DSP modules

  • Because of the ADC readout problem (see 5-Jan-2007 entry), the BSM23E DSP module (sn005) has been swapped for the designated XrayBSM unit (sn002).

January 15, 2007 (R.Holtzapple, E. Tanke): CESR-c beamsize measurements, both species in the machine.

  • Present CESR-C conditions are that we have current in the following 24 bunches:
    • T1B3,T1B4
    • T2B3,T2B4,T2B5
    • T3B3,T3B4,T3B5
    • T4B3,T4B4,T4B5
    • T5B3,T5B4,T5B5
    • T6B3,T6B4,T6B5
    • T7B3,T7B4,T7B5
    • T8B3,T8B4,T8B5
    • T9B5
  • A typical run is currently ~ 33 minutes. 10k turn (raw) measurements made during one and the same run using T1B1 as pedestal bunch:
    • Beginning of run: file 943 (electrons) and file 944 (positrons)
    • Middle of run: file 945 (electrons) and file 946 (positrons)
    • End of run: file 947 (electrons) and file 948 (positrons)

February 2007

February 06, 2007 (E.Tanke): BSM23E files up to and including #1005 are now on the server

  • BSM23E files up to and including #1005 are now on the server in the /nfs/cesr/temp/ts101/BSM23E_data directory

March 2007

March 13, 2007 (G.Codner, R.Holtzapple, J.Kern, M.Palmer, L.Schachter, E.Tanke): MS ECLOUD CESR-c with witness bunches

  • Timing scan at PMT HV=-500V confirms Tglob=500 as optimum (file 1044)
  • Positrons nominally loaded to 0.75 mA/bunch, with a leading train of 10 bunches, followed by one or more witness bunches. Files are for 8k turns of raw data. Only the witness bunches are shown in the table below (all measurements include bunches 1 through 10):
    • #1046: B42
    • #1047: B38,42
    • #1048: B34,38,42
    • #1049: B30,34,38,42
    • #1050: B28,30,34,38,42
    • #1051: B26,28,30,34,38,42
    • #1052: B24,26,28,30,34,38,42 (but B25 was the bunch that got filled in stead of 24)
    • #1053: B22,25,26,28,30,34,38,42
    • #1054: B20,22,25,26,28,30,34,38,42
    • #1055: B19,20,22,25,26,28,30,34,38,42 (but no current in B19)
    • #1056: B19,20,22,25,26,28,30,34,38,42
    • #1057: B18,19,20,22,25,26,28,30,34,38,42 (right after this measurement a beam loss occurred)
    • #1058: B17,18,19,20,22,25,26,28,30,34,38,42
    • #1059: B16,17,18,19,20,22,25,26,28,30,34,38,42
    • #1060: B15,16,17,18,19,20,22,25,26,28,30,34,38,42
    • #1061: B14,15,16,17,18,19,20,22,25,26,28,30,34,38,42
    • #1062: B13,14,15,16,17,18,19,20,22,25,26,28,30,34,38,42
    • #1063: B12,13,14,15,16,17,18,19,20,22,25,26,28,30,34,38,42
    • #1064: B11,12,13,14,15,16,17,18,19,20,22,25,26,28,30,34,38,42

April 2007

April 02, 2007 (G.Codner, R.Holtzapple, J.Kern, M.Palmer, E.Tanke): MS ECLOUD CESR-c with witness bunches

  • Timing scan at PMT HV=-500V confirms Tglob=500 as optimum (file 1097)
  • Positrons nominally loaded to 0.75 mA/bunch, with a leading train of 10 bunches, followed by one or more witness bunches. Files are for 1k turns of raw data. Only the witness bunches are shown in the table below (all measurements include bunches 1 through 10):
    • #1100: B42
    • #1101: B38,42
    • #1102: B34,38,42
    • #1103: B30,34,38,42
    • #1104: B26,30,34,38,42
    • #1105: B22,26,30,34,38
    • #1106: B20,22,26,30,34
    • #1107: B19,20,22,26,30
    • #1108: B18,19,20,22,26
    • #1109: B17,18,19,20,22
    • #1110: B16,17,18,19,20
    • #1111: B15,16,17,18,19
    • #1112: B14,15,16,17,18
    • #1113: B13,14,15,16,17
    • #1114: B12,13,14,15,16
    • #1115: B12,13,14,15,16,17,18,19,20
    • #1116: B11,12,13,14,15,16,17,18,19,20
    • #1117: B11,12,13,14,15,16,17,18,19,20 (HW trigger 1, but shutter overcurrent)
    • #1118: B11,12,13,14,15,16,17,18,19,20 (HW trigger 1, 8k turns, but shutter overcurrent)

April 05, 2007 (E.Tanke): New trigger handling for HW trigger 1 (bit 0)

  • When in triggered mode, the BSM23E (and BSM23W) unit will sit and wait for this trigger. As soon as it sees the trigger the shutter will be opened after which there is a fixed delay of 70k turns before data taking starts (one needs to wait for the shutter to open and for mechanical vibrations to settle). Note that this is the way these units now function ONLY for taking raw data and ONLY for trigger bit 0. This can be expanded as needed.

April 11, 2007 : /nfs server directory restructuring

  • The /nfs server directory has been restructured (see this link). BSM23E files that have been stored under /nfs/cesr/temp/ts101/BSM23E_data before, can now be found under /nfs/acc/temp/ts101/BSM23E_data

June 2007

June 05, 2007 (M.Palmer) : Lens moved 3.75" downstream

  • In order to change the magnification of the light seen by the PMT, the lens has been moved 3.75" downstream (i.e. closer to the PMT). The 2 mirrors in the dogleg have been slightly moved such as to horizontally center the light on the PMT.

June 08, 2007 (E.Tanke) : Modified FE of BSM sn005 in view of operation with 4 ns bunch spacing

  • In view of operation with 4 ns bunch spacing, BSM sn005 has been modified. This unit is currently in the DAQlab (the spare unit foreseen for the Xray BSM experiments is in the tunnel). A characterization of the response of this BSM can be found here and the corresponding detailed measurement data can be found here.

  • All 32 channels have been tested to be operational.

June 08, 2007 (J.Dobbins) : Details of BSM Front End modification

  • On each channel 3 capacitors were removed. See attached schematic showing a typical channel. C19, C20, C77 were removed.

June 12, 2007 (E.Tanke) : BSM sn005 installed at 23E; calibration measurements performed

  • BSM sn005 (unit modified for 4 ns bunch spacing operation) has been installed at 23E, BSM sn002 is now in the DAQlab.
  • Measurements in view of calibrating BSM23E using the CCD linear array data obtained with VSIZE as reference; VB56 is used to change the position of the light on the PMT and on the CCD (VB56=0 is nominal setting; VB56 was taken through a hysteresis loop):
VB56 (cu) CCD file # BSM23E file #
-60 73229 1204
-40 73319 1205
-20 73407 1206
0 73504 1207
20 73551 1208
40 73707 1210
60 73752 1211
80 73835 1212
100 74014 1213
Topic revision: r16 - 18 Jun 2007, EugeneTanke
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding CLASSE Wiki? Send feedback